OpenStudy (tay0913):

the measure of two of the sides of an equilateral triangle are 3x+15 in. and 7x-5 in. What is the measure of the third side in inches?

OpenStudy (anonymous):

3x+15=7x-5 3x-7x=-5-15 -4x=-20 4x=20 x=4*20 (haha) x=80

OpenStudy (decarr432):

Welcome to open study and @zactowbes please do not just give the answer help them through it

OpenStudy (tay0913):

Thank u @decarr432 but I got 30 and still don't understand how ^^ @zactowbes got there answer

OpenStudy (tay0913):

and this is geometry . just adding that

OpenStudy (decarr432):

Thats okay as long as its math you can post it in this subject

OpenStudy (tay0913):

k. but I still need help solving this

OpenStudy (decarr432):

I suck at this I am sorry but I know some one who may be able to help @mathmale

jhonyy9 (jhonyy9):

@zactowbes so what you have wrote above is wrong than 4x=20 so x not is equal 80 how you get it ?

jhonyy9 (jhonyy9):

@Decarr432 can you solve it for medal ?

OpenStudy (decarr432):

Not likely sir I apologize but geometry is my worst subject but I can try

OpenStudy (decarr432):

Wouldn't x have to be equal to 5 because 3(5)+15=30 and 7(5)-5=30

OpenStudy (decarr432):

@jhonyy9

jhonyy9 (jhonyy9):

@Decarr432 first of all what mean exactly ,,equilateral triangle" ?

OpenStudy (decarr432):

A triangle in which all three sides are equal right?

jhonyy9 (jhonyy9):

yes - so than you know that a side has length of 3x+15 and the second 7x-5 so can you calculi x from these two ?

OpenStudy (decarr432):

X is 5 3(5)+15=30 and 7(5)-5=30

jhonyy9 (jhonyy9):

so than what will be the length of 3rd side ?

OpenStudy (decarr432):

30

jhonyy9 (jhonyy9):

exactly - nice work -

OpenStudy (decarr432):

Thank you!

jhonyy9 (jhonyy9):

np was my pleasure good luck bye

OpenStudy (decarr432):

@Tay0913 Did you see how I got that?

OpenStudy (tay0913):

Yes. Thank u

OpenStudy (tay0913):

Both of u guys @decarr432 & @jhonyy9

OpenStudy (decarr432):

If you need help just tag me I'll do my best make sure to close this btw! :D

OpenStudy (tay0913):

Ok

OpenStudy (tay0913):

Both of u guys @decarr432 & @jhonyy9